A jet plane lands with a speed of 112 m/s and can accelerate at a maximum rate of −5.50 m/s2 as it comes to rest. (a) From the instant the plane touches the runway, what is the minimum time needed before it can come to rest? s (b) Can this plane land on a small tropical island airport where the runway is 0.800 km long? Yes No
Answer #1

a) Applying v = u + at

for final speed, v = 0

0 = 112 - 5.50t

t = 20.36 sec


b) Distance needed to stop,

v^2 - u^2 = 2ad

0^2 - 112^2 =2(-5.50)d

d = 1140.36 m Or 1.14 km

so 0.8km is not sufficient.

Ans - No